# A Finer View of the Parameterized Landscape of Labeled Graph Contractions

### v1

@@ 3.1.3 Updating Dynamic Programming Tables
($\Leftarrow$) Assume the conditions are met. Condition (i) must hold since $u,v\in V(H)$, meaning they must be the roots of their witness sets, which cannot merge due to a $G$-edge. By the premise, a valid partial solution $P^{\prime}$ exists for $G_{t^{\prime}}$. We construct $P$ for $G_{t}$ by adding $(u,v)$. Since the functions $\sigma_{p},\tau_{p}$ and $\mathcal{A}_{p}$ are unchanged (condition (i) and (ii)), the only effect is to satisfy the $H$-requirement between $\tau(u)$ and $\tau(v)$. Condition (iii) correctly formalizes this satisfaction, updating an unsat requirement to sat in $\mathcal{R}$. Thus $P$ is a valid partial solution for $G_{t}$.
($\Rightarrow$) If $\mathbf{d}[S,t]=\text{true}$, a valid $P$ exists, derived from a valid $P^{\prime}$ at $t^{\prime}$. Since $u,v\in V(H)$, the edge $(u,v)$ cannot cause a witness set merge, so the functions must remain the same, proving condition (i). The edge is a $G$-edge, not a pseudo-adjacency, so $\mathcal{A}$ remains $\mathcal{A}^{\prime}$, proving condition (ii). Its only effect is on the $H$-contraction requirement, which means if $(u,v,\textsf{unsat})$ was in $\mathcal{R}^{\prime}$, it must be resolved to sat in $\mathcal{R}$. If it was already sat or inv, $\mathcal{R}$ must be the same as $\mathcal{R}^{\prime}$. This verifies condition (iii). ∎
Lemma 3.7.
Let $t$ be an introduce edge node where edge $(u,v)$ is introduced, with $u\in V(H)$ and $v\notin V(H)$. A signature $S$ is valid at node $t$ if and only if there exists a signature $S^{\prime}\in I_{e}(S)$ such that $\mathbf{d}[S^{\prime},t^{\prime}]=\texttt{True}$, and one of the following holds:
  1. (i)
    Condition C2.1 holds and $\mathcal{R}=\mathcal{R}^{\prime}$ and $\mathcal{A}=\mathcal{A}^{\prime}\cup\{(u,v)\}$.
  2. (ii)
    Condition C2.2 holds and $\mathcal{R}=\mathcal{R}^{\prime}$ and $\mathcal{A}=\mathcal{A}^{\prime}$.
Proof.
($\Leftarrow$) Assume one of the conditions holds. A valid $S^{\prime}$ exists by the premise.
  1. 1.
    If (i) holds, the functions are preserved (C2.1), meaning no witness set merge occurs. The edge $(u,v)$ is thus recorded as a new pseudo-adjacency in $\mathcal{A}$, forming a valid $P$ for $G_{t}$.
  2. 2.
    If (ii) holds, the functions change according to C2.2, meaning the witness set containing $v$ merges into $u$’s set. The edge $(u,v)$ is absorbed into the new witness structure and does not become a pseudo-adjacency, so $\mathcal{A}$ remains $\mathcal{A}^{\prime}$. This forms a valid $P$ for $G_{t}$.
($\Rightarrow$) If $\mathbf{d}[S,t]=\text{true}$, a valid $P$ exists. Since $u\in V(H)$ and $v\notin V(H)$, $u$ must be a requester and $v$ cannot be the root of a witness set. The new edge $(u,v)$ must either form a pseudo-adjacency (case (i)) or cause $v$’s set to be absorbed by $u$’s (case (ii)). Case (i) requires function stability (C2.1). Case (ii) requires the witness set merge, which corresponds to the $\sigma$ and $\tau$ changes in C2.2. These two cases are exhaustive and mutually exclusive for a valid solution. ∎
Lemma 3.8.
Let $t$ be an introduce edge node where edge $(u,v)$ is introduced, with $u\notin V(H)$ and $v\in V(H)$. A signature $S$ is valid at node $t$ if and only if there exists a signature $S^{\prime}\in I_{e}(S)$ such that $\mathbf{d}[S^{\prime},t^{\prime}]=\texttt{True}$, and one of the following holds:
  1. (i)
    Condition C2.1 holds and $\mathcal{A}=\mathcal{A}^{\prime}\cup\{(u,v)\}$.
  2. (ii)
    Condition C2.3 holds and $\mathcal{R}=\mathcal{R}^{\prime}$ and $\mathcal{A}=\mathcal{A}^{\prime}$.
Proof.
This is symmetric to Lemma 3.7, with $u$ and $v$ roles swapped. Case (i) is for pseudo-adjacency (no merge), and Case (ii) is for $u$’s set merging into $v$’s set (C2.3). The proof follows the same logic. ∎
Lemma 3.9.
Let $t$ be an introduce edge node where edge $(u,v)$ is introduced, with $u\notin V(H)$ and $v\notin V(H)$. A signature $S$ is valid at node $t$ if and only if there exists a signature $S^{\prime}\in I_{e}(S)$ such that $\mathbf{d}[S^{\prime},t^{\prime}]=\texttt{True}$, and one of the following holds:
  1. (i)
    Condition C2.1 holds and $\mathcal{A}=\mathcal{A}^{\prime}\cup\{(u,v)\}$.
  2. (ii)
    Condition C2.2 holds and $\mathcal{R}=\mathcal{R}^{\prime}$ and $\mathcal{A}=\mathcal{A}^{\prime}$.
  3. (iii)
    Condition C2.3 holds and $\mathcal{R}=\mathcal{R}^{\prime}$ and $\mathcal{A}=\mathcal{A}^{\prime}$.
Proof.
This case combines the possibilities from the previous two lemmas. Since neither $u$ nor $v$ are $H$-vertices, the edge $(u,v)$ can either: (i) be a pseudo-adjacency (functions unchanged, C2.1), or (ii) cause $v$’s set to merge into $u$’s (C2.2), or (iii) cause $u$’s set to merge into $v$’s (C2.3). In the merge cases ((ii) and (iii)), the edge is absorbed into the witness structure, leaving $\mathcal{A}$ unchanged. The proof follows the same construction and existence logic as the previous cases. ∎
Join Node
Let $t$ be a join node in the tree decomposition $T$, with children $t_{1}$ and $t_{2}$, such that $X_{t}=X_{t_{1}}=X_{t_{2}}$. Let $S=(\sigma,\tau,\mathcal{R},\mathcal{A})$ be a signature at node $t$, and let $S_{1}=(\sigma_{1},\tau_{1},\mathcal{R}_{1},\mathcal{A}_{1})$ and $S_{2}=(\sigma_{2},\tau_{2},\mathcal{R}_{2},\mathcal{A}_{2})$ be signatures at nodes $t_{1}$ and $t_{2}$ respectively.
Definition 3.15.
We say that $S$, $S_{1}$, and $S_{2}$ are join-node compatible if the following conditions hold:
  • (C1) For every $y\in X_{t}$, we have $\sigma(y)=\sigma_{1}(y)=\sigma_{2}(y)$ and $\tau(y)=\tau_{1}(y)=\tau_{2}(y)$.
  • (C2) The signature sets are combined by union: $\mathcal{R}=\mathcal{R}_{1}\cup\mathcal{R}_{2}$ and $\mathcal{A}=\mathcal{A}_{1}\cup\mathcal{A}_{2}$.
Let $J(S)$ denote the set of all pairs of signatures $(S_{1},S_{2})$ at $t_{1}$ and $t_{2}$ respectively, that satisfy (C1) and (C2) with $S$.
Lemma 3.10.
Let $t$ be a join node with children $t_{1}$ and $t_{2}$, and let $X_{t}=X_{t_{1}}=X_{t_{2}}$. Let $S=(\sigma,\tau,\mathcal{R},\mathcal{A})$ be a signature at node $t$. Then, $\mathbf{d}[S,t]=\texttt{True}$ if and only if there exists a pair of signatures $(S_{1},S_{2})\in J(S)$, where $S_{1}=(\sigma_{1},\tau_{1},\mathcal{R}_{1},\mathcal{A}_{1})$ is a signature at $t_{1}$ and $S_{2}=(\sigma_{2},\tau_{2},\mathcal{R}_{2},\mathcal{A}_{2})$ is a signature at $t_{2}$, such that:
  1. (i)
    $\mathbf{d}[S_{1},t_{1}]=\texttt{True}$ and $\mathbf{d}[S_{2},t_{2}]=\texttt{True}$.
  2. (ii)
    The sets $\mathcal{A}$ and $\mathcal{R}$ are formed by combining the corresponding sets from $S_{1}$ and $S_{2}$ according to the following rules for every pair $(u,v)\in X_{t}\times X_{t}$:
    • $\mathcal{A}=\mathcal{A}_{1}\cup\mathcal{A}_{2}$.
    • $\mathcal{R}$ is the union of all combined requester entries, $\mathcal{R}=\bigcup_{u,v\in X_{t}}\mathcal{R}(u,v)$, where $\mathcal{R}(u,v)$ is determined by the $\oplus$ and $\otimes$ operators:
      • Requester Merge ($\oplus$): For $(u,v,\ell_{1})\in\mathcal{R}_{1}$ and $(u,v,\ell_{2})\in\mathcal{R}_{2}$:
[…]
      • Pseudo-Adjacency Satisfies ($\otimes$): For $(u,v,\ell_{1})\in\mathcal{R}_{1}$ and $(u,v)\in\mathcal{A}_{2}$:
        $\ell_{1}\otimes\mathcal{A}_{2}=\begin{cases}\textsf{sat}&\text{if }\ell_{1}=\textsf{unsat}\\ \ell_{1}&\text{if }\ell_{1}\in\{\textsf{sat},\textsf{inv}\}\end{cases}$
        (A symmetric rule applies for $\mathcal{R}_{2}\otimes\mathcal{A}_{1}$.)
      The resulting entry for $(u,v)$ in $\mathcal{R}$ is $\ell_{1}\oplus\ell_{2}$ (if both exist), or the result of the $\otimes$ operations, prioritizing sat over unsat and unsat over non-existence, and rejecting inv.
Proof.
$(\Leftarrow)$ Assume conditions (i) and (ii) hold. By condition (i), there exist valid partial solutions $P_{1}=(\sigma_{p1},\tau_{p1},\mathcal{R}_{p1},\mathcal{A}_{p1})$ for $G_{t_{1}}$ and $P_{2}=(\sigma_{p2},\tau_{p2},\mathcal{R}_{p2},\mathcal{A}_{p2})$ for $G_{t_{2}}$. We construct a new partial solution $P=(\sigma_{p},\tau_{p},\mathcal{R}_{p},\mathcal{A}_{p})$ for $G_{t}=G_{t_{1}}\cup G_{t_{2}}$. The functions $\sigma_{p}$ and $\tau_{p}$ are defined piece-wise by the union of $\sigma_{p1},\tau_{p1}$ and $\sigma_{p2},\tau_{p2}$. Since $X_{t}=X_{t_{1}}=X_{t_{2}}$, condition (C1) ensures the functions agree on $X_{t}$, guaranteeing that the union is a valid structure for the witness sets. The set of edges $E(G_{t})=E(G_{t_{1}})\cup E(G_{t_{2}})$.
  • Pseudo-Adjacency ($\mathcal{A}$): The pseudo-adjacencies in $P$ are exactly the union of those in $P_{1}$ and $P_{2}$, as $G_{t_{1}}$ and $G_{t_{2}}$ share no edges. Condition (ii) enforces $\mathcal{A}=\mathcal{A}_{1}\cup\mathcal{A}_{2}$.
  • Requester Set ($\mathcal{R}$): The requester set tracks $H$-requirements between the originators of vertices in $X_{t}$. A requirement is satisfied in $G_{t}$ if it is satisfied in $G_{t_{1}}$ or $G_{t_{2}}$, or if the necessary edge is provided by a pseudo-adjacency in the other subgraph.
    • The $\oplus$ operator correctly determines the status based on the union of $G_{t_{1}}$ and $G_{t_{2}}$: an unsat requirement becomes sat if the edge is found in the other subgraph (represented by an sat in the other $\mathcal{R}$ set).
    • The $\otimes$ operator correctly models the satisfaction of an unsat requirement in one subgraph ($\mathcal{R}_{1}$ or $\mathcal{R}_{2}$) by a pseudo-adjacency recorded in the other subgraph ($\mathcal{A}_{2}$ or $\mathcal{A}_{1}$).
    Condition (ii) correctly computes the final status for $\mathcal{R}$ ensuring all constraints are properly inherited and resolved.
Since $P_{1}$ and $P_{2}$ are valid, and $P$ is constructed consistently with the union rules, $P$ is a valid partial solution for $G_{t}$. The signature of $P$ is $S$, thus $\mathbf{d}[S,t]=\texttt{True}$.
$(\Rightarrow)$ Assume $\mathbf{d}[S,t]=\text{true}$. There exists a valid partial solution $P$ for $G_{t}$ whose signature is $S$. By the definition of $G_{t}$ and tree decomposition, $P$ must be the result of combining two valid partial solutions $P_{1}$ for $G_{t_{1}}$ and $P_{2}$ for $G_{t_{2}}$.
  • Let $S_{1}$ and $S_{2}$ be the signatures of $P_{1}$ and $P_{2}$, respectively. Since $P_{1}$ and $P_{2}$ are valid, condition (i) must hold.
  • Since $P$ is a valid union of $P_{1}$ and $P_{2}$ and $X_{t}$ is the intersection, the function values must agree on $X_{t}$, which verifies (C1) and ensures $(S_{1},S_{2})\in J(S)$.
  • The pseudo-adjacency set $\mathcal{A}$ in $S$ must be the union of the pseudo-adjacencies recorded in $S_{1}$ and $S_{2}$ as the edges are disjoint, verifying $\mathcal{A}=\mathcal{A}_{1}\cup\mathcal{A}_{2}$.
  • The final status of any $H$-requirement $(u,v)$ in $\mathcal{R}$ must be the result of combining the status from $S_{1}$ and $S_{2}$, accounting for satisfaction via: (1) internal edges in $G_{t_{1}}$, (2) internal edges in $G_{t_{2}}$, or (3) pseudo-adjacencies in the opposite subgraph. This requires the exact combination defined by the $\oplus$ and $\otimes$ operators in condition (ii) to hold.
Since a valid $P$ exists, a pair $(S_{1},S_{2})$ satisfying all specified conditions must exist. ∎
3.1.4 Putting it all together.
We are now in a position to present the proof of Theorem 1 which states that Labeled Contraction admits an algorithm with running time $2^{\mathcal{O}(\textsf{tw}^{2})}\cdot|V(G)|^{\mathcal{O}(1)}$.
Proof.
The proof relies on the Dynamic Programming (DP) approach defined by the signatures $S=(\sigma,\tau,\mathcal{R},\mathcal{A})$ and the update laws (Lemmas 3.2 through 3.10). The overall running time is determined by the number of nodes in the tree decomposition, the size of the DP state space, and the time required for each transition.
At any node $t$ of the tree decomposition, the size of the bag $|X_{t}|$ is at most $\textsf{tw}+1$. The number of distinct signatures $S=(\sigma,\tau,\mathcal{R},\mathcal{A})$ stored at node $t$ is bounded by considering each component: The Parent Function $\sigma:X_{t}\to X_{t}\cup\{\star\}$ and the Originator Function $\tau:X_{t}\to V(H)\cup\{\star\}$ each contribute $2^{\mathcal{O}(\textsf{tw}\log\textsf{tw})}$ states. For $\tau$, this is because only $\mathcal{O}(\textsf{tw})$ vertices in $V(H)$ are relevant to the bag $X_{t}$. The Requester Set $\mathcal{R}$ encodes a status from $\{\textsf{sat},\textsf{unsat},\textsf{inv}\}$ for $\mathcal{O}(\textsf{tw}^{2})$ pairs, resulting in $3^{\mathcal{O}(\textsf{tw}^{2})}=2^{\mathcal{O}(\textsf{tw}^{2})}$ choices. The Pseudo-Adjacency Set $\mathcal{A}$ is binary for $\mathcal{O}(\textsf{tw}^{2})$ pairs, contributing $2^{\mathcal{O}(\textsf{tw}^{2})}$ choices. Multiplying these components yields the total number of distinct signatures:
$\text{Total States}=2^{\mathcal{O}(\textsf{tw}\log\textsf{tw})}\cdot 2^{\mathcal{O}(\textsf{tw}\log\textsf{tw})}\cdot 2^{\mathcal{O}(\textsf{tw}^{2})}\cdot 2^{\mathcal{O}(\textsf{tw}^{2})}=2^{\mathcal{O}(\textsf{tw}^{2})}$
The DP algorithm processes each of the $\mathcal{O}(|V(G)|)$ nodes in the tree decomposition. At Leaf, Introduce, and Forget Nodes, the transition involves iterating over the $2^{\mathcal{O}(\textsf{tw}^{2})}$ signatures of the single child node, with the compatibility checks and updates for each taking time polynomial in the bag size, $\mathcal{O}(\textsf{tw}^{\mathcal{O}(1)})$. At Join Nodes, the transition requires checking all pairs of compatible signatures $(S_{1},S_{2})$ from the two child nodes. Since there are $2^{\mathcal{O}(\textsf{tw}^{2})}$ signatures at each child, there are at most $\left(2^{\mathcal{O}(\textsf{tw}^{2})}\right)^{2}=2^{\mathcal{O}(\textsf{tw}^{2})}$ pairs to examine, and the combining operations ($\oplus$ and $\otimes$) for each pair take time $\mathcal{O}(\textsf{tw}^{\mathcal{O}(1)})$. Consequently, the time required to process any single node is bounded by $2^{\mathcal{O}(\textsf{tw}^{2})}\cdot\textsf{tw}^{\mathcal{O}(1)}=2^{\mathcal{O}(\textsf{tw}^{2})}$.
The overall running time of the algorithm is the number of nodes in the decomposition multiplied by the time per node:
$\text{Time}=\mathcal{O}(|V(G)|)\cdot 2^{\mathcal{O}(\textsf{tw}^{2})}$
This is bounded by $2^{\mathcal{O}(\textsf{tw}^{2})}\cdot|V(G)|^{\mathcal{O}(1)}$. This concludes the proof of the algorithmic result mentioned in Theorem 1. ∎
@@ 3.2 Conditional Lower Bound
In this section, we demonstrate that unless (ETH) fails, Labeled Contraction does not admit an algorithm running in time $2^{o(\textsf{tw}^{2})}\cdot|V(G)|^{\mathcal{O}(1)}$, where tw is the treewidth of the input graph $G$. To obtain the lower bound, we reduce from a special case of Vertex Cover on sub-cubic graphs, defined in [3] as follows:
@@ 5.1 Conditional Lower Bound
We now proceed with the proof of Theorem 3.
@@ 5.2 Algorithmic Result
We first present proof of Theorem 4 establishing an alternate $\mathsf{FPT}$ algorithm and later mention its consequences. Recall that the theorem states that the Labeled Contraction problem admits an algorithm with running time $(\delta(H)+1)^{k}\cdot|V(G)|^{\mathcal{O}(1)}$, where $k=|V(G)\setminus V(H)|$.
@@ 6 Brute-force Algorithm and its Optimality
We now prove Theorem 5. We restate it for reader’s conveience.
@@
Now we present the conditional lower bound, demonstrating that the algorithm is essentially optimal unless the Exponential Time Hypothesis (ETH) fails.

### v2

@@ 3.1 Algorithmic Result
(Algorithm in Theorem 1) The algorithm employs dynamic programming over a nice tree decomposition of $G\cup H$, maintaining states indexed by signatures $S=(\mathcal{P},\tau,\mathcal{A}_{t},\mathcal{R}_{t})$. For a bag $X_{t}$ of size at most $\textsf{tw}+1$, the total number of distinct signatures $|\mathcal{S}|$ is bounded by combining the state space components. The number of partitions $\mathcal{P}$ of $X_{t}$ is bounded by $2^{\mathcal{O}(\textsf{tw}\log\textsf{tw})}$. The status function $\tau$, which maps each part to one of $\mathcal{O}(\textsf{tw})$ possible representatives or status markers ($\uparrow,\downarrow$), yields at most $(\textsf{tw}+3)^{|\mathcal{P}|}=2^{\mathcal{O}(\textsf{tw}\log\textsf{tw})}$ possibilities. Since the number of pairs of parts is $\mathcal{O}(\textsf{tw}^{2})$, the total number of combinations for the adjacency sets $\mathcal{A}_{t}$ and $\mathcal{R}_{t}$ is $2^{\mathcal{O}(\textsf{tw}^{2})}$. Combining these, the total number of signatures per node is $|\mathcal{S}|\leq 2^{\mathcal{O}(\textsf{tw}^{2})}$. Each DP transition at Leaf, Introduce, and Forget nodes iterates over $\mathcal{O}(|\mathcal{S}|)$ states of the child, performing checks in polynomial time in tw. At Join nodes, pairs of states are combined, resulting in $\mathcal{O}(|\mathcal{S}|^{2})=2^{\mathcal{O}(\textsf{tw}^{2})}$ combined states, merged in polynomial time. Since the tree decomposition has $\mathcal{O}(|V(G)|)$ nodes, the overall running time is $\mathcal{O}(|V(G)|)\cdot 2^{\mathcal{O}(\textsf{tw}^{2})}=2^{\mathcal{O}(\textsf{tw}^{2})}\cdot|V(G)|^{\mathcal{O}(1)}$. This complexity proves the correctness of the algorithm mentioned in Theorem 1. ∎
@@ 3.2 Conditional Lower Bound
In this section, we demonstrate that unless the ETH fails, Labeled Contraction does not admit an algorithm running in time $2^{o(\textsf{tw}^{2})}\cdot|V(G)|^{\mathcal{O}(1)}$, where tw is the treewidth of the input graph $G$. To obtain the lower bound, we reduce from a special case of Vertex Cover on sub-cubic graphs, defined in [3] as follows:
@@ 5.1 Conditional Lower Bound
We now proceed with the proof of Theorem 3 which we restate here.
See 3
@@ 5.2 Algorithmic Result
We first present proof of Theorem 4 establishing an alternate $\mathsf{FPT}$ algorithm and later mention its consequences. We restate the theorem.
See 4
Proof.
@@ 6 Brute-force Algorithm and its Optimality
We now prove Theorem 5. We restate it for reader’s convenience.
@@
6.1 Brute-force Algorithm
We prove the following lemma.
@@ 6.1 Brute-force Algorithm
6.2 Conditional Lower Bound
In this subsection, we present the conditional lower bound, demonstrating that the algorithm is essentially optimal unless the Exponential Time Hypothesis (ETH) fails.